# Continuous-time Data-driven Barrier Certificate Synthesis

### v1

@@ 4 Certificate Synthesis
2:     Set $k\leftarrow 1$triangleright Initialise iteration index
3:     Set $\mathcal{C}\leftarrow\varnothing$triangleright Initialise compression set
4:     Fix $L_{1}<L_{0}$ with $|L_{1}-L_{0}|>\eta$triangleright $\eta$ is any fixed tolerance
@@
6:         $g\leftarrow\nabla_{\theta}l^{s}(\theta)$triangleright Gradient of loss function
7:         $\theta\leftarrow\theta-\alpha g$triangleright Step in the direction of sample-independent gradient
@@
10:         $\mathcal{M}\leftarrow\{\tilde{\xi}\in\mathcal{D}\colon L(\theta,\tilde{\xi})\geq\max_{\tilde{\xi}\in\mathcal{C}}L(\theta,\tilde{\xi})\}$triangleright Find maximal samples with loss greater than compression set loss
11:         $\overline{g}_{\mathcal{M}}\leftarrow\{\nabla_{\theta}L(\theta,\tilde{\xi})\}_{\tilde{\xi}\in\mathcal{M}}$triangleright Subgradients of loss function for $\tilde{\xi}\in\mathcal{M}$
12:         $\overline{\xi}_{\mathcal{C}}\in\operatorname*{arg\,max}_{\tilde{\xi}\in\mathcal{C}}L(\theta,\tilde{\xi})$triangleright Find a sample with maximum loss from $\mathcal{C}$
13:         $\overline{g}_{\mathcal{C}}\leftarrow\nabla_{\theta}L(\theta,\overline{\xi}_{\mathcal{C}})$triangleright Approximate subgradient of loss function for $\tilde{\xi}=\overline{\xi}_{\mathcal{C}}$
@@
16:              $\theta\leftarrow\theta-\alpha\overline{g}$triangleright Step in the direction of misaligned subgradient
17:              $\mathcal{C}\leftarrow\mathcal{C}\cup\{\overline{\xi}\}$triangleright Update compression set with sample corresponding to $\overline{g}$
@@
19:              $\theta\leftarrow\theta-\alpha\overline{g}_{\mathcal{C}}$triangleright Step in the direction of approximate subgradient
@@
21:         $L_{k}\leftarrow\min\left\{L_{k-1},\max_{\tilde{\xi}\in\mathcal{D}}L(\theta,\tilde{\xi})\right\}$triangleright Update “running” loss value
22:         $k\leftarrow k+1$triangleright Update iteration index
@@
Figure 1: Graphical Representation of Algorithm 1.
4.1 Certificate and Compression Set Computation
We seek to minimize a loss function that encodes the barrier certificate conditions, with respect to the neural network parameters. To this end, for a $\tilde{\xi}\in\tilde{\Xi}$ and parameter vector $\theta$, let
@@ 4.1 Certificate and Compression Set Computation
$l^{s}(V_{\theta})$ :$\vcentcolon=\frac{1}{|\mathcal{X}_{I}|}\sum_{x\in\mathcal{X}_{I}}\max\{0,V_{\theta}(x)\}$
$+\frac{1}{|\mathcal{X}_{U}|}\sum_{x\in\mathcal{X}_{U}}\max\{0,-V_{\theta}(x)\}.$
$l^{\Delta}(\theta,\tilde{\xi})$ :$\vcentcolon=-\frac{1}{T}\Big{(}\inf_{x\in\mathcal{X}_{U}}B(x)-\sup_{x\in\mathcal{X}_{I}}B(x)\Big{)}$
$+\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}.$
@@
To instantiate these functions we consider a discrete set of grid-points on each sub-domain: $\mathcal{X}_{\overline{I}}$ is the set of points in the domain but outside the goal region, and $\mathcal{X}_{U}$ is the set of points in the unsafe set. These points are generated densely enough across the domain of interest, and hence offer an accurate approximation. Since these samples do not require access to the dynamics, we consider them separate to the sample-set $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
Consider the first summation in the sample-independent loss, if $B(x)<0$ then $\max\{0,B_{\theta}(x)\}=0$, i.e., no loss is incurred, implying satisfaction of (3). Under a similar reasoning, the other integral accounts for (4), respectively.
We impose the next mild assumption, a sufficient condition for termination of our algorithm.
Assumption 5 (Minimizers’ Existence).
For any $\{\tilde{\xi}\}_{i=1}^{N}$, and any non-empty $\mathcal{D}\subseteq\{\tilde{\xi}\}_{i=1}^{N}$, the set of minimizers of $\max_{\tilde{\xi}\in\mathcal{D}}L(\theta,\tilde{\xi}),$ is non-empty.
Note that, for a sufficiently expressive neural network, we can find a certificate $B$ which satisfies the state constraints and hence has a sample-independent loss of zero.
@@
Algorithm 1 provides an inexact subgradient methodology to minimize the loss function, and to iteratively construct a compression set $\mathcal{C}$ (initially empty; see step $3$). We explain the main steps of this algorithm with reference to Figure 1. After an arbitrary parameter initialization, we follow the subgradient associated with the current sample in $\mathcal{C}$ (“blue” dot labeled by $3$). At this point, this step becomes inexact, as there would exist another sample resulting in a higher loss (“green” point labeled by $2$). Such a sample is in $\mathcal{M}$, step $9$ of Algorithm 1. However, the algorithm does not “jump” to the green point, as the condition in step $13$ of the algorithm is not yet satisfied. As such the algorithm performs inexact subgradient descent steps up to point $5$; this is the first instance where the condition in step $13$ is satisfied (there exists another constraint with opposite slope) and hence the algorithm “jumps” to a point with higher loss and subgradient of opposite sign. The “jumps” serve as an exploration step to investigate the non-convex landscape, while their number corresponds to the cardinality of the returned compression set. Such a procedure can be thought of as a constructive procedure for the general framework presented recently in [16] to construct compression sets.
In some cases, the parameter returned by Algorithm 1 may result in a value of the loss function greater than $-d$, and hence mean we are unable to verify safety. To achieve a lower loss, we make use of a sample-and-discarding procedure [6, 20], and introduce Algorithm 2 as an outer loop around Algorithm 1. This procedure leads to a lower loss, however the samples that are discarded have to be added to the compression set. Theorems 3 and 4 then apply, but with the compression set returned by Algorithm 2 instead.
To terminate our algorithm we require knowledge of $d$, but cannot calculate $d$ until we find $\theta^{\star}$. To resolve this, we propose two different approaches.
@@
    At every iteration $j$ calculate $d_{j}$ using the current best parameters $\theta_{j}$, terminate when $\max_{i}\left[L(\theta_{j},\tilde{\xi}^{i})\right]<d_{j}$.
@@
    Choose a parameter set $\Theta$, and take the supremum across the set to find an upper bound on $\mathcal{L}_{B},\mathcal{M}_{B}$, use these to calculate $d$.
To determine Lipschitz constants for neural networks we refer the reader to [4, 9, 22].
@@
3:Set $\mathcal{D}\leftarrow\{\tilde{\xi}^{i}\}_{i=1}^{N}$triangleright Initialise “running” samples
@@
5:     $\theta,\mathcal{C}\leftarrow$ $\mathcal{A}(\theta,\mathcal{D})$triangleright Call Algorithm 1
6:     $\widetilde{\mathcal{C}}\leftarrow\widetilde{\mathcal{C}}\cup\mathcal{C}$triangleright Update $\widetilde{\mathcal{C}}$
7:     $\mathcal{D}\leftarrow\mathcal{D}\setminus\widetilde{\mathcal{C}}$triangleright Discard $\widetilde{\mathcal{C}}$ from $\mathcal{D}$
@@
5 Numerical Results
We consider constructing a safety certificate for the nonlinear, two-dimensional jet engine model as considered in [15],
$\dot{x}_{1}(t)=-x_{2}(t)-\frac{3}{2}x_{1}^{2}(t)-\frac{1}{2}x_{1}^{3}(t),\;\;\dot{x}_{2}(t)=x_{1}(t).$ (17)
In Figure 3 we provide a graphical representation of the dynamics, sub-domains under study, the 00-level set produced by our certificate, and the level sets calculated by the methods in [15] (one lower bounding the unsafe set, the other upper bounding the initial set). We used $5$ independent repetitions (each with different multi-samples) of $1,000$ sampled trajectories and $367$ seconds of computation time (standard deviation $139$s), to obtain $\varepsilon=0.01492$ (standard deviation $0.00140$) with confidence to $0.99$. The methodology of [15] required $257149$ state pair samples and $5123$ seconds (standard deviation $449$s) of computation time to compute a barrier certificate with the same confidence (however, this holds deterministically). We estimate the Lipschitz constants using the methods in [23]. Figure 3 contains a 3D plot of the certificate.
Figure 2: Phase plane plot, initial and unsafe set for (17). The zero-level set for our certificate is dashed; level sets that bound the initial and unsafe sets (i.e. $\gamma$- and $\lambda$- level sets) in [15] are dotted.
@@ 5 Numerical Results
Beyond these numerical results, we briefly discuss the theoretical differences between our approaches. The results in [15] offer a guarantee that, with a certain confidence, the safety property is always satisfied, in contrast to Theorem 4 where we provide such guarantees in probability (up to a quantifiable risk level $\varepsilon$). However, these “always” guarantees, albeit very useful, come with some challenges. Firstly, they are not applicable when part of the initial set is unsafe. Secondly, they implicitly require some knowledge of the underlying probability distribution. Finally, they are bound to an exponential growth with the dimension of the state space.
@@
$\dot{x}_{3}(t)$ $=0.01\sqrt{|x_{1}(t)|},$
$\dot{x}_{4}(t)$ $=-x_{1}(t)-x_{2}(t)^{2}+\sin(x_{4}(t)).$
We calculate that the approach of [15] requires at least $10^{19}$ samples to return a confidence lower bounded by at most $10^{-30}$, which is not practically useful. In contrast, our techniques with only $100$ samples obtain a risk level of $\varepsilon=0.21450$ (standard deviation $0.00910$), confidence $1-10^{-5}$.
6 Conclusion
We have proposed a method for synthesis of neural-network certificates for continuous-time dynamical systems, based only on a finite number of trajectories from a system. Our numerical experiments demonstrate the efficacy of our methods on a number of examples, involving comparison with related methodologies in the literature. Current work concentrates towards extending our analysis to controlled systems, thus co-designing a controller and a certificate at the same time.
@@ 6 Conclusion
1 Proof of Theorem 4
We aim at finding a bound on the discretisation gap $L(\theta,\xi)-L(\theta,\tilde{\xi})$ so that, for sufficiently small loss evaluated on the time-discretised approximations $L(\theta,\tilde{\xi})$, we also achieve a negative loss on the continuous trajectories $L(\theta,\xi)$.
$L(\theta,\xi)-L(\theta,\tilde{\xi})=l^{\Delta}(\theta,\xi)-l^{\Delta}(\theta,\tilde{\xi})$
$=\max_{x\in\xi}\evaluated{\frac{dB}{dt}}_{x}-\max_{k=1,\dots,M}\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}.$ (19)
Replacing the first maximisation with one between time instances, and exchanging the order of the $\max$ operators, (19) is equal to
$\max_{k=1,\dots,M}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}$ (20)
@@ 1 Proof of Theorem 4
$\leq\max_{k=1,\dots,M}\Bigg{[}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}-\frac{B(x(t_{k}))-B(x(t_{k-1}))}{t_{k}-t_{k-1}}\Bigg{]}.$ (21)
We can now replace the difference term with an integral, so that (21) becomes equivalent to
$\max_{k=1,\dots,M}\frac{\int_{t_{k-1}}^{t_{k}}\max_{t\in[t_{k-1},t_{k}]}\evaluated{\frac{dB}{dt}}_{x(t)}-\evaluated{\frac{dB}{dt}}_{x(\tau)}~{}\mathrm{d}\tau}{t_{k}-t_{k-1}}.$

### v2

@@ IV Certificate Synthesis
2:  Set $k\leftarrow 1$ $\triangleright$ Initialise iteration index
3:  Set $\mathcal{C}\leftarrow\varnothing$ $\triangleright$ Initialise compression set
4:  Fix $L_{1}<L_{0}$ with $|L_{1}-L_{0}|>\eta$ $\triangleright$ $\eta$ is any fixed tolerance
@@
6:   $g\leftarrow\nabla_{\theta}l^{s}(\theta)$ $\triangleright$ Gradient of loss function
7:   $\theta\leftarrow\theta-\alpha g$ $\triangleright$ Step in the direction of sample-independent gradient
@@
[…]
11:   $\overline{g}_{\mathcal{M}}\leftarrow\{\nabla_{\theta}L(\theta,\tilde{\xi})\}_{\tilde{\xi}\in\mathcal{M}} \triangleright$ Subgradients of loss function for $\tilde{\xi}\in\mathcal{M}$
[…]
13:   $\overline{g}_{\mathcal{C}}\leftarrow\nabla_{\theta}L(\theta,\overline{\xi}_{\mathcal{C}})$ $\triangleright$ Approximate subgradient of loss function for $\tilde{\xi}=\overline{\xi}_{\mathcal{C}}$
@@
16:     $\theta\leftarrow\theta-\alpha\overline{g}$ $\triangleright$ Step in the direction of misaligned subgradient
17:     $\mathcal{C}\leftarrow\mathcal{C}\cup\{\overline{\xi}\}$ $\triangleright$ Update compression set with sample corresponding to $\overline{g}$
@@
19:     $\theta\leftarrow\theta-\alpha\overline{g}_{\mathcal{C}}$ $\triangleright$ Step in the direction of approximate subgradient
@@
[…]
22:   $k\leftarrow k+1$ $\triangleright$ Update iteration index
@@
IV-A Certificate Synthesis Algorithm
We seek to minimise a loss function that encodes the barrier certificate conditions, with respect to the neural network parameters. For $\tilde{\xi}\in\tilde{\Xi}$ and parameter vector $\theta$, let
@@ IV-A Certificate Synthesis Algorithm
$l^{s}(\theta)$ $\vcentcolon=\frac{1}{|\mathcal{X}_{I}|}\sum_{x\in\mathcal{X}_{I}}\max\{0,B_{\theta}(x)\}$
$+\frac{1}{|\mathcal{X}_{U}|}\sum_{x\in\mathcal{X}_{U}}\max\{0,\delta-B_{\theta}(x)\}.$
[…]
$+\max_{k=1,\dots,M}\frac{B_{\theta}(x(t_{k}))-B_{\theta}(x(t_{k-1}))}{t_{k}-t_{k-1}},$
@@
To instantiate these functions we consider a discrete set of $N_{\mathrm{states}}$ grid-points on each sub-domain: $\mathcal{X}_{I}$ is the set of points in the initial set, and $\mathcal{X}_{U}$ is the set of points in the unsafe set. These points are generated densely enough across the domain of interest, and hence offer an accurate approximation. Since they do not require access to the dynamics, we consider them separately from $\{\tilde{\xi}^{i}\}_{i=1}^{N}$.
[…]
The following mild assumption ensures a minimiser of the loss functions exists and hence our algorithm terminates.
Assumption 3 (Minimisers’ Existence)
[…]
@@
Algorithm 1 provides an inexact subgradient methodology to minimise the loss function, and to iteratively construct a compression set $\mathcal{C}$ (initially empty; see step $3$). We explain the main steps of this algorithm with reference to Figure 1, where each sample gives rise to a concave triangular constraint. After an arbitrary initialization, we follow the subgradient associated with the worst case sample and add it to the compression set $\mathcal{C}$ (step 14141515, point $M_{1}$). When iterates get to point $M_{2}$, the subgradient step becomes inexact, as for the same parameter there exists a sample resulting in a higher loss (see asterisk). Such a sample is in $\mathcal{M}$, step $9$ of Algorithm 1. However, the algorithm does not “jump”, as the condition in step $13$ of the algorithm is not yet satisfied. Graphically, this is since the $M_{2}$ and the red asterisk are on a side of the respective constraint with the same slope. As such the algorithm performs inexact subgradient descent steps up to point $M_{3}$; this is the first instance where the condition in step $13$ is satisfied (there exists another constraint with opposite slope) and hence the algorithm “jumps” to a point with higher loss and subgradient of opposite sign. This procedure is then repeated as shown in the figure, with the red line indicating the iterates’ path. The “jumps” serve as an exploration step to investigate the non-convex landscape, while their number corresponds to the cardinality of the returned compression set. This can be thought of as a constructive procedure for the general framework presented in [17] to construct compression sets.
The computational complexity of one iteration of the main loop is $\mathcal{O}((M+N_{\mathrm{states}})\cdot N\cdot N_{\mathrm{layers}}\cdot N_{\mathrm{neurons}}^{3})$. We can upper bound the number of iterations by $\frac{L_{0}-L^{\star}}{\eta}$, where $L^{\star}$ is the value of the loss associated with one of the minimisers. To terminate our algorithm we require knowledge of $d$ (which depends on $\theta^{\star}$). To resolve this, we propose two approaches.
@@
    At every iteration $j$ calculate $d_{j}$ using the current best parameters $\theta_{j}$, terminate if $\max_{i}\left[L(\theta_{j},\tilde{\xi}^{i})\right]<-d_{j}$.
@@
    Choose a parameter set, and consider the supremum across that set to determine an upper bound on $\mathcal{L}_{B},\mathcal{M}_{B}$. Use these upper bounds to calculate $d$.
The second option is likely to be conservative but removes the need for calculating Lipschitz constants in the loop, hence is computationally more efficient. To determine Lipschitz constants for neural networks we refer to [18, 19].
[…]
Figure 1: Graphical illustration of Algorithm 1.
IV-B Sample Discarding and Compression Computation
Due to the tightening introduced, to minimise the loss function, we would like its value to be lower than or equal to d-d. However, in some cases, the parameter returned by Algorithm 1 may result in a loss value greater than d-d, thus preventing us from verifying safety. To ensure the loss is below that threshold, we employ a sampling-and-discarding procedure [20, 21], and introduce Algorithm 2 as an outer loop around Algorithm 1. Algorithm 2 iterates as long the desired loss value is not met (step $4$), and progressively discards samples (step $7$). The samples that are discarded are the ones identified as compression by Algorithm 1 (step $6$), as removing these is bound to improve the solution.
@@ IV-B Sample Discarding and Compression Computation
3:Set $\mathcal{D}\leftarrow\{\tilde{\xi}^{i}\}_{i=1}^{N}$ $\triangleright$ Initialise “running” samples
@@
5:  $\theta,\mathcal{C}\leftarrow$ $\mathcal{A}(\theta,\mathcal{D})$ $\triangleright$ Call Algorithm 1
6:  $\widetilde{\mathcal{C}}\leftarrow\widetilde{\mathcal{C}}\cup\mathcal{C}$ $\triangleright$ Update $\widetilde{\mathcal{C}}$
7:  $\mathcal{D}\leftarrow\mathcal{D}\setminus\widetilde{\mathcal{C}} \triangleright$ Discard $\widetilde{\mathcal{C}}$ from $\mathcal{D}$
@@
The samples that are discarded need to be added to the compression set (step $6$); intuitively this is the case as if we repeat the procedure using only the compression samples we need to follow the same solution path. A direct byproduct of Algorithms 1 and 2 is that they calculate a compression set, whose cardinality $C_{N}$ is in turn used in Theorem 2 to provide the desired probabilistic guarantees. The cardinality of that set corresponds to the number of “jumps” in Algorithm 1 plus the number of discarded samples in Algorithm 2. This cardinality is not necessarily greater for higher dimensional problems, but is rather case dependent and depends on the complexity of the problem. For example, a problem where some trajectories approach or even enter the unsafe set presents a more challenging synthesis problem than one where trajectories all move in the opposite direction to the unsafe set, thus we expect the former to have a larger compression set even if the problem is smaller in dimension.
V Numerical Results
We consider constructing a safety certificate for the nonlinear, two-dimensional jet engine model as considered in [2], The codebase is available at https://github.com/lukearcus/fossil_scenario
$\dot{x}_{1}(t)=-x_{2}(t)-\frac{3}{2}x_{1}^{2}(t)-\frac{1}{2}x_{1}^{3}(t),\;\;\dot{x}_{2}(t)=x_{1}(t),$ (15)
using a time horizon $T=5$. Figure 3 provides a graphical representation of the dynamics, subdomains under study, the 0-level set produced by our certificate, and the level sets calculated by the methods in [2] (one lower bounding the unsafe set, the other upper bounding the initial set). We used $5$ independent repetitions (each with different multi-samples) of $1,000$ sampled trajectories and $367$ seconds of computation time (standard deviation $139$s), to obtain $\varepsilon=0.01492$ (standard deviation $0.00140$) with confidence $0.99$. The methodology of [2] required $257149$ state pair samples and $5123$ seconds (standard deviation $449$s) of computation time to compute a barrier certificate with the same confidence (however, this holds deterministically). We estimate the Lipschitz constants using the methodology in [22], noting that convergence is guaranteed asymptotically. Figure 3 contains a 3D plot of the certificate.
Figure 2: Phase plane plot, initial and unsafe set for (15). The zero-level set for our certificate is dashed; level sets that bound the initial and unsafe sets in [2] are dotted.
@@ V Numerical Results
Beyond these numerical results, we briefly discuss the theoretical differences between our approach and [2]. The results in [2] offer a guarantee that, with a certain confidence, the safety property is always satisfied, in contrast to Theorem 2 where we provide such guarantees in probability (up to a quantifiable risk level $\varepsilon$). However, these “always” guarantees, although very useful, come with some challenges. Firstly, they are not applicable when part of the initial set is unsafe, whereas we can still bound the probability of a new trajectory being safe even if some of the sampled trajectories were unsafe. Second, they implicitly require knowledge of the underlying probability distribution to instantiate their confidence bound and scale exponentially with respect to the state space dimension.
@@
[…]
$\dot{x}_{4}(t)$ $=-x_{1}(t)-x_{2}(t)^{2}+\sin(x_{4}(t)),$
using $T=4$. We applied also the approach of [2] which, with $10^{19}$ samples, returned a confidence $10^{-30}$, an uninformative result as it is close to zero. With only $100$ samples our techniques obtain a risk level $\varepsilon=0.21450$ (standard deviation $0.00910$), confidence $1-10^{-5}$ (close to one).
VI Conclusion
We have proposed a method for synthesis of neural-network certificates for continuous-time dynamical systems, based only on a finite number of trajectories from a system. Our numerical experiments demonstrate the efficacy of our methods on a number of examples, involving comparison with related methodologies in the literature. Current work concentrates towards extending to controlled systems, thus co-designing a controller and a certificate at the same time.
@@ VI Conclusion
I Proof of Theorem 2
[…]
Replace the first maximisation with one between time instances, and exchange the order of the $\max$ operators,
[…]
@@ I Proof of Theorem 2
[…]
We can now replace the difference term with an integral,
[…]